OpenStudy (anonymous):

The highway mileage m in miles per gallon for a compact car is approximated by m(s) = -0.025s^2 + 2.45s - 30, where s is the speed in miles per hour. What is the maximum mileage for this compact car to the nearest tenth of a mile per gallon? __________ mi/gal.

OpenStudy (anonymous):

x value of the vertex= -a/2b inverse of that -b/2a = -2.45/2(-0.025)= 49 plug in 49 for S

OpenStudy (anonymous):

Do I square s first?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, I got 30.025 but it said it was wrong. Is that right? Or did I do something wrong?

OpenStudy (anonymous):

it just 30 cuz you're only worried about to the 1/10 of the mile

OpenStudy (anonymous):

Oh okay, thank you! So do you know how I could figure out the speed that results from this mileage?

OpenStudy (anonymous):

And 30 isn't right either :(

OpenStudy (anonymous):

do u have idea about derivation.

OpenStudy (anonymous):

No, I don't think so.

OpenStudy (anonymous):

can u convert it into completed square form

OpenStudy (anonymous):

-0.025s^2 + 2.45s - 3

OpenStudy (anonymous):

m(s) = -0.025s^2 + 2.45s - 30 =-0.025(s^2-98s)-30 =-0.025{s^2 - 2 * s * 49 + 49^2 -49^2)-30 =-0.025{(s-49)^2 -2401}-30 =-0.025(s-49)^2 + 60.025-30 =-0.025(s-49)^2+30.025

OpenStudy (anonymous):

for all real values of s,